If a rowing team takes 2 hours to travel 16 miles with the current but takes 4 miles to travel the same distance against the current what is their rate of speed in still water?
Thank you!

Answer by mananth(16949)

Rowing speed =x mph
current speed =y mph
against current x-y 2.00 hours
withcurrent x+y 2.00 hours

Distance = same= 2000 miles
t=d/r
4 / ( x - y )= 2.00
2 ( x - y ) = 4.00
2 x -2 y = 4 ....................1
16 / ( x + y )= 2.00
2.00 ( x + y ) = 4
2.00 x + 2.00 y = 16 ...............2
Multiply (1) by 1.00
Multiply (2) by 1.00
we get
2 x + -2 y = 4
2 x + 2 y = 16
4 x = 20
/ 4
x = 5 mph speed of rowing

plug value of x in (1)
2 x -2 y = 4
10 -2 y = 4
-2 y = 4 -10
-2 y = -6
y = 3 mph current speed